OpenStudy (anonymous):

John shot 67 baskets that were worth 2 or 3 points each. He made a total of 160 points. How many of each type of basket did John make? Hint* Set up a system of equations.

OpenStudy (gw2011):

Let x = 2 point baskets Let y = 3 point baskets 2x + 3y = 160 x + y = 67 Solving the above equations by eliminating x: 2x + 3y = 160 -2x - 2y = -134 y = 26 x = 41 Therefore, 26 three-point baskets and 41 two-point baskets were made